# DeepScholar-Bench: A Live Benchmark and Automated Evaluation for Generative Research Synthesis

### v1

@@ 5. Experimental Results
Experimental Setup. For each benchmarked method, we control the retrieval corpus by allowing each system to access the Web only through the ArXiv API (noa, [n.d.]a). We additionally avoid possible information leakage during search by filtering out any search results that were published after the query paper’s publication date. We report results using GPT-4.1-2025-04-14 (noa, [n.d.]i) as the judge for Nugget Coverage, and a GPT-4o-2024-08-06 (noa, [n.d.]i) judge for Organization, Relevance Rate, Reference Coverage, Citation Precision and Claim Coverage. We report the Organization score as a win rate including ties, we report the strict all score for Nugget Coverage, and we report Claim Coverage with a window size of . For all Retrieval Quality metrics, we consider the retrieved set of each given report as the set of any valid ArXiv links found within the report. To measure Document Importance, we use the OpenAlex (noa, [n.d.]j) API to recover citation information. For each metric, we report an average over all reports.
5.1. Baselines
We briefly overview all of the baseline systems we evaluate, and we provide further details in the Appendix Section 8.2.
5.1.1. Open-source Research Systems
@@
We evaluate three state-of-the-art open-source systems, DeepResearcher (Zheng et al., 2025), STORM (Shao et al., 2024) and OpenScholar (Asai et al., 2024). For each, we run these systems using the Llama-4-Scout-17B-16E-Instruct model (noa, 2025c), which we serve with 4 A100 GPUs using vLLM (Kwon et al., 2023).
@@
DeepResearcher (Zheng et al., 2025) leverages trained agents to navigate, browse and synthesize information from the web. To train an agent, this work uses end-to-end reinforcement learning and trains Qwen2.5-7B-Instruct (Qwen et al., 2025). In our benchmarks, we evaluate DeepResearcher using both the released, trained model from the authors, and using Llama-4-Scout-17B-16E-Instruct model (noa, 2025c) as the core LLM. We report the better performing baseline of these two, which we find in our experiments to be the Llama-4-Scout-17B-16E-Instruct backbone.
@@
STORM (Shao et al., 2024) studies the problem of how to apply LLMs to write grounded, organized long-form articles (e.g., Wikipedia articles) from scratch. The system involves a pre-writing stage that discovers diverse research perspectives on a topic by stimulating conversations between multiple agents and leveraging web documents.
@@
OpenScholar (Asai et al., 2024) builds a specialized retrieval-augmented LLM system for literature synthesis and scientific queries. This method includes a trained retriever from the pre-indexed peS2o (Soldaini et al., 2024) corpus, consisting of 45 million open-access academic papers up until October 2024, as an initial retrieval source before using web search. In our experiments, we benchmark the system using this pre-indexed corpus and limit web search to the ArXiv API.
5.1.2. Search AI’s
@@
We evaluate the following models: Llama-4-Scout-17B-16E-Instruct (noa, 2025c), GPT-4.1-2025-04-14 (noa, [n.d.]i), o3-2025-04-16 (noa, [n.d.]i), Claude-opus-4-20250514 (noa, [n.d.]f), and Gemini-2.5-pro (noa, [n.d.]d). We augment each with search capabilities to ArXiv (noa, [n.d.]a), and use the popular ODS framework (Alzubi et al., 2025) to allow the LLM to make tool calls to the search API.
5.1.3. Commercial Systems.
@@
We focus our evalution of commercial generative research synthesis systems on OpenAI’s o3-deep-research (noa, [n.d.]i), which provides a public API allowing for our evaluation.
5.1.4. DeepScholar-base
@@
Similar to our evaluation of search AI’s we evaluate DeepScholar-base with the following models: Llama-4-Scout-17B-16E-Instruct (noa, 2025c), GPT-4.1-2025-04-14 (noa, [n.d.]i), o3-2025-04-16 (noa, [n.d.]i), Claude-opus-4-20250514 (noa, [n.d.]f), and Gemini-2.5-pro (noa, [n.d.]d). For each of these baselines, we also use the same or a weaker model, either Llama-4 or GPT-4.1, to perform semantic filtering and top-k operators. We limit the method to two round of search, each with at most 2 queries.
5.2. Main Results
Table 2 provides detailed summary of each method’s performance scores on all metrics across three key dimensions, knowledge synthesis, retrieval quality and verifiability. We also provide metadata statistics characterizing the generated reports of each benchmarked method in Table 5, as well as statistics related to our evaluation metrics in Table 6. Overall, our evaluation demonstrates two key findings, which we discuss in detail below: first, existing generative research synthesis systems demonstrate significant headway for improvement, and second, DeepScholar-base provides a strong baseline for generative research synthesis.
| Knowledge Synthesis | Retrieval Quality | Verifiability | |||||
| Org. | Nug. Cov. | Rel. Rate | Ref Cov. | Doc Imp. | Cite-P | Claim Cov () | |
@@ 5.2. Main Results
| Human-written Exemplars | .500 | 1.000 | .585 | 1.000 | 1.000 | .278111The verifiability metrics we use in our evaluation likely significantly under-estimate the actual verifiability of human writing. This is because the metrics we measure, Citation Precision and Claim Coverage, require us to verify entailment relations between claims, within the report, and snippets from the cited reference. For each LLM-based system, we are able to track the precise snippet and context from each cited source, which are directly fed as context to the LLM as it generates the report with citations. On the other hand, for the human-written exemplars, we lack gold labels pointing to the precise snippet of text that each reference refers to. Our measurements for the human-written exemplars instead rely on the title and abstract of each cited source when testing for entailment relations. | .205111The verifiability metrics we use in our evaluation likely significantly under-estimate the actual verifiability of human writing. This is because the metrics we measure, Citation Precision and Claim Coverage, require us to verify entailment relations between claims, within the report, and snippets from the cited reference. For each LLM-based system, we are able to track the precise snippet and context from each cited source, which are directly fed as context to the LLM as it generates the report with citations. On the other hand, for the human-written exemplars, we lack gold labels pointing to the precise snippet of text that each reference refers to. Our measurements for the human-written exemplars instead rely on the title and abstract of each cited source when testing for entailment relations. |
@@
| DeepResearcher (Llama-4) | .206 | .230 | .385 | .047 | .008 | .312 | .396 |
| STORM (Llama-4) | .119 | .183 | .218 | .003 | .006 | .238 | .586 |
| OpenScholar (Llama-4) | .309 | .278 | .017 | .008 | .013 | .010 | .138 |
| Search AI’s | |||||||
| Search AI (Llama-4-Scout) | .151 | .193 | .445 | .060 | .009 | .316 | .368 |
| Search AI (GPT-4.1) | .556 | .265 | .490 | .050 | .009 | .498 | .470 |
| Search AI (o3) | .849 | .348 | .610 | .165 | .026 | .425 | .495 |
| Search AI (Claude) | .698 | .307 | .583 | .131 | .008 | .701 | .760 |
| Search AI (Gemini) | .706 | .277 | .583 | .061 | .010 | .415 | .398 |
@@
| OpenAI DeepResearch | .857 | .392 | .629 | .187 | .124 | .399 | .138 |
| DeepScholar Baseline | |||||||
| DeepScholar-base (Llama-4) | .254 | .262 | .421 | .103 | .008 | .648 | .826 |
| DeepScholar-base (GPT-4.1) | .825 | .407 | .608 | .162 | .007 | .652 | .636 |
| DeepScholar-base (GPT-4.1, o3) | .857 | .405 | .659 | .162 | .008 | .617 | .614 |
| DeepScholar-base (GPT-4.1, Claude) | .786 | .370 | .586 | .167 | .007 | .936 | .817 |
| DeepScholar-base (GPT-4.1, Gemini) | .762 | .332 | .602 | .181 | .006 | .851 | .865 |
5.2.1. Generative Research Synthesis Systems Demonstrate Large Room for Improvement.
From Table 2, we see that no method is able to achieve a score greater than across all metrics. Moreover, on several key metrics, including nugget coverage, reference coverage and document importance, each evaluated method’s performance remains well below . This reflects the inherent difficulty of the generative research task provided by DeepScholar-bench, which requires systems to navigate the live web, reasoning about relevance and importance of documents to perform retrieve sources and then surface key information into a coherent report that answers the query.
We now analyze each evaluated dimension, comparing performance of the open-source research systems, search AI’s and commercial systems to the human-written exemplars. On knowledge synthesis, we see that OpenAI DeepResearch offers the best performance compared to all other prior methods on both Organization, with a score of , and Nugget Coverage, with a score of . OpenAI DeepResearch, as well as the o3, Claude and Gemini search AI’s achieve relatively high Organization scores compared to human-written exemplars. However, on Nugget Coverage all prior methods scores below . This demonstrates that while existing systems, especially those using state-of-the-art models, can generate well-organized and coherent summaries, they still struggle to extract and surface key facts, a crucial capability for synthesis tasks.
Turning our attention to the retrieval quality performance of prior methods, we once again find significant room for improvement. Once again, OpenAI DeepResearch offers the strongest performance among the other benchmarked prior methods on Relevance Rate, Reference Coverage and Document Importance, but still far from saturates performance. While it’s Relevance Rate shows strong performance, exceeding that of the human exemplars with a score of , it’s Reference Coverage and Document Importance scores remain exceedingly low: and respectively. This demonstrates that while state-of-the-art generative research synthesis systems are capable of retrieving relevant sources, they still struggle to find a comprehensive set of notable sources and fall short compared to the ability of human experts.
Lastly, we analyze the verifiability performance of prior methods, we see that OpenAI DeepResearch is outperformed on both Citation Precision and Claim Coverage by the search AI’s with GPT4.1, o3, Claude and Gemini models. The Claude search AI offers the highest Citation Precision, a score of and Claim Coverage, a score of . Meanwhile, OpenAI’s DeepResearch as well as the all other prior methods are unable to achieve a Citation Precision score beyond and a Claim Coverage score beyond . We also note that the human-written exemplars appear to exhibit rather low Citation Precision and Claim Coverage scores, however these scores are not comparable to the metric measured for the LLM-based systems since our method for measuring verifiability metrics likely under-estimate the actual verifiability of human writing11footnotemark: 1. Overall, we see that prior LLM-based systems exhibit significant headroom for improvement.
| Knowledge Synthesis | Retrieval Quality | Verifiability | |||||
| Org | Nug. Cov. | Rel. Rate | Ref Cov. | Doc Imp. | Cite-P | Claim Cov () | |
| DeepScholar-base (GPT-4.1, Claude) | |||||||
| arxiv.org Retrieval | .786 | .370 | .586 | .167 | .007 | .936 | .817 |
| parallel.ai Retrieval | .865 | .444 | .675 | .160 | .017 | .846 | .781 |
| taviliy.com Retrieval | .929 | .327 | .550 | .070 | .015 | .711 | .578 |
| Oracle Retrieval (arxiv.org) | .782 | .487 | .686 | 1.000 | 1.000 | .955 | .899 |
| Oracle Retrieval (All) | .778 | .528 | .680 | 1.000 | .822 | .941 | .828 |
| DeepScholar-base (Llama-4) | |||||||
| arxiv.org Retrieval | .254 | .262 | .421 | .103 | .008 | .648 | .826 |
| parallel.ai Retrieval | .246 | .265 | .559 | .114 | .015 | .223 | .543 |
| taviliy.com Retrieval | .111 | .229 | .532 | .030 | .016 | .442 | .676 |
| Oracle Retrieval (arxiv.org) | .202 | .316 | .681 | 1.000 | 1.000 | .658 | .868 |
| Oracle Retrieval (All) | .198 | .350 | .693 | 1.000 | .822 | .796 | .890 |
@@
| Retrieval Relevance Score | Graded Relevance (0/1/2) | |
| Reference Coverage | Reference Importance (Not Imp./ Imp.) |
5.2.2. DeepScholar-base Provides a Strong Baseline for Generative Research Synthesis.
We compare the performance of DeepScholar-base to the commercial OpenAI DeepResearch system, search AI’s and open-source research systems, finding that DeepScholar-base provides competitive performance against each group of prior methods, offering a strong baseline for generative research synthesis.
First, we see that DeepScholar-base (GPT-4.1, o3) is competitive with OpenAI DeepResearch, achieving a similar or higher Organization, Nugget Coverage, Relevance Rate, Reference Coverage, Citation Precision and Claim Coverage scores. Notably, DeepScholar-base offers significantly higher verifiability than OpenAI DeepResearch, with higher Citation Precision and higher Claim Coverage. However, DeepScholar-base’s document importance scores still remain especially low compared to OpenAI DeepResearch, representing significant room for improvement.
Next, we compare the performance of DeepScholar-base to the search-AI’s, finding that for each evaluated model, DeepScholar-base consistently offers improved performance compared to the corresponding search AI. Specifically, averaged across all 5 baselines with different models for the Search AI’s and DeepScholar-base method, DeepScholar-base offers higher Organization, higher Nugget Coverage, higher Relevance Rate, higher Reference Coverage higher Citation Precision, higher Citation Recall.
Lastly, we compare DeepScholar-base (Llama-4) to the open-source research systems, all run with the Llama-4 model as well. We see that the prior open-source research systems exhibit trade-offs among the Knowledge Synthesis, Retrieval Quality and Verifiability dimensions. Specifically, OpenScholar achieves the highest Knowledge Synthesis scores, on both Organization and Nugget Coverage, DeepResearcher achieves the highest Relevance Rate and Reference Coverage on Retrieval Quality, with all systems attaining only very low Document Importance scores, and on Verifiability, DeepResearcher offers the highest Citation-Precision while STORM offers the highest Claim Coverage. In comparison, DeepScholar-base offers strong performance across each dimension. Specifically, Compared to the best-performing prior open-source methods for each metric, DeepScholar-base offers competitive Knowledge Synthesis performance, higher Relevance Rates and higher Reference Coverage for retrieval, and higher Citation Precision and higher Claim Coverage scores for verifiability.
Overall, the strong relative performance of DeepScholar-base likely reflects the efficiency of the data-processing semantic operators (Patel et al., 2025) that DeepScholar-base uses to perform LLM-based filtering, ranking and summarization of sources to generate its report. Notably, DeepScholar-base still demonstrates significant room for improvement and far from saturates DeepScholar-Bench, especially on key Knowledge Synthesis and Retrieval Quality metrics, including Nugget Coverage, Reference Coverage, Document Importance, which represent opportunities for future work.
5.3. Understanding Opportunities for Improvement.
| Report Length | Citations | ||||
| Chars | Words | Sentences | # Unique Refs | # Inline Citations | |
| Human-written Exemplars | |||||
| Human-written Exemplars | 4381 | 497 | 28 | 23 | 27 |
| Open Source Research Systems | |||||
| DeepResearcher (Llama-4) | 2573 | 319 | 35 | 8 | 7 |
| STORM (Llama-4) | 2766 | 381 | 31 | 18 | 21 |
| OpenScholar (Llama-4) | 3513 | 483 | 26 | 9 | 19 |
| Search AI’s | |||||
| Search AI (Llama-4-Scout) | 1968 | 258 | 16 | 9 | 5 |
| Search AI (GPT-4.1) | 3168 | 404 | 16 | 10 | 61 |
| Search AI (o3) | 3844 | 501 | 24 | 11 | 16 |
| Search AI (Claude) | 3977 | 499 | 27 | 13 | 8 |
| Search AI (Gemini) | 2810 | 395 | 19 | 6 | 8 |
| Commercial Systems | |||||
| OpenAI DeepResearch | 6577 | 864 | 74 | 17 | 6 |
| DeepScholar Baseline | |||||
| DeepScholar-base (Llama-4) | 3864 | 402 | 58 | 21 | 19 |
| DeepScholar-base (GPT-4.1) | 14470 | 1492 | 167 | 19 | 56 |
| DeepScholar-base (GPT-4.1, o3) | 5905 | 642 | 69 | 16 | 20 |
| DeepScholar-base (GPT-4.1, Claude) | 12287 | 1332 | 136 | 23 | 35 |
| DeepScholar-base (GPT-4.1, Gemini) | 6118 | 663 | 81 | 22 | 28 |
| Avg. value over human-written exemplars | Relevant Metric | |
| # Important References from ArXiv.org | 11.47 | Ref. Cov. |
In order to further analyze performance and existing opportunities for improvement, we conduct an ablation study, where we consider different retrievers as well as two oracle retriever settings. Table 3 shows these results for two evaluated DeepScholar-base methods, DeepScholar-base (GPT-4.1, Claude) and DeepScholar-base (Llama-4). The table show the performance of either using three different retrieval APIs, including arxiv.org, the default used in our main results, parallel.ai and tavily.com. In addition the table shows to oracle settings for either DeepScholar-base method: the Oracle Retrieval (arxiv.org) setting, provides the system with the ArXiv references from the human-written exemplars labeled as ”Important” following our methodology for evaluating Reference Coverage. The Oracle Retrieval (All) setting, provides the system with any from the human-written exemplars labeled as ”Important” following the same methodology, including both references form ArXiv and once that are found elsewhere.
Overall,the results shown in Table 3 demonstrate that the performance limitations of existing systems for generative research synthesis lie in both their retrieval capabilities to find and select high-quality references sets, as well as their synthesis abilities to surface key facts and extract insights given the retrieved documents.
First, we see that given either oracle retrieval setting, the DeepScholar-base (GPT-4.1, Claude) method nearly saturates performance on Retrieval Quality and Verifiability metrics, whereas the same method using the arxiv.org, parallel.ai or taviliy.com retrievers obtain lower scores on each of these metrics. This finding demonstrates a significant opportunity to improve performance of generative research systems through improvements to the retrieval method. Specifically, existing systems struggle to find a diverse and holistic set of notable sources, reflected by their especially low Reference Coverage and Document Importance scores.
Additionally, we also see that oracle retrieval settings for either DeepScholar-base method attain higher Nugget Coverage, improving the score by up to compared to the rspective arxiv.org, parallel.ai or tavily.ai retreival settings. However, we note that the oracle retreival methods still far from saturate Nugget Coverage, with the DeepScholar-base (GPT-4.1, Claude) Oracle Retrieval (All) attaining a modest score of .. This demonstrates that even with high retrieval quality, existing LLM systems still struggle to effectively surface important facts and synthesize important insights.
5.4. Human Agreement Study

### v2

@@ A.2 Overview of Baselines and Experimental Setup
We report results using GPT-4.1-2025-04-14 [34] as the judge for Nugget Coverage, and a GPT-4o-2024-08-06 [34] judge for Organization, Relevance Rate, Reference Coverage, Citation Precision and Claim Coverage. We report the Organization score as a win rate including ties, we report the strict all score for Nugget Coverage, and we report Claim Coverage with a window size of . For all Retrieval Quality metrics, we consider the retrieved set of each given report as the set of any valid ArXiv links found within the report. To measure Document Importance, we use the OpenAlex [35] API to recover citation information. For each metric, we report an average over all reports.
A.2.1 Open-source Research Systems
@@
We evaluate three state-of-the-art open-source systems, DeepResearcher [65], STORM [43] and OpenScholar [6]. For each, we run these systems using the Llama-4-Scout-17B-16E-Instruct model [30], which we serve with 4 A100 GPUs using vLLM [22].
@@
DeepResearcher [65] leverages trained agents to navigate, browse and synthesize information from the web. To train an agent, this work uses end-to-end reinforcement learning and trains Qwen2.5-7B-Instruct [40]. In our benchmarks, we evaluate DeepResearcher using both the released, trained model from the authors, and using Llama-4-Scout-17B-16E-Instruct model [30] as the core LLM. We report the better performing baseline of these two, which we find in our experiments to be the Llama-4-Scout-17B-16E-Instruct backbone.
@@
STORM [43] studies the problem of how to apply LLMs to write grounded, organized long-form articles (e.g., Wikipedia articles) from scratch. The system involves a pre-writing stage that discovers diverse research perspectives on a topic by stimulating conversations between multiple agents and leveraging web documents.
@@
OpenScholar [6] builds a specialized retrieval-augmented LLM system for literature synthesis and scientific queries. This method includes a trained retriever from the pre-indexed peS2o [44] corpus, consisting of 45 million open-access academic papers up until October 2024, as an initial retrieval source before using web search. In our experiments, we benchmark the system using this pre-indexed corpus and limit web search to the ArXiv API.
A.2.2 Search Agents
@@
We evaluate the following models: Llama-4-Scout-17B-16E-Instruct [30], GPT-4.1-2025-04-14 [34], o3-2025-04-16 [34], Claude-opus-4-20250514 [3], and Gemini-2.5-pro [14]. We augment each with search capabilities to ArXiv [5], and use the popular ODS framework [1] to allow the LLM to make tool calls to the search API.
A.2.3 Commercial Systems.
@@
We focus our evaluation of commercial generative research synthesis systems on OpenAI’s o3-deep-research [34], which provides a public API allowing for our evaluation.
A.2.4 DeepScholar-ref
Similar to our evaluation of search agent we evaluate DeepScholar-ref with the following models:
@@ A.2.4 DeepScholar-ref
Llama-4-Scout-17B-16E-Instruct [30], GPT-4.1-2025-04-14 [34], o3-2025-04-16 [34], Claude-opus-4-20250514 [3], and Gemini-2.5-pro [14]. For each of these baselines, we also use the same or a weaker model, either Llama-4 or GPT-4.1, to perform semantic filtering and top-k operators. We limit the method to two rounds of search, each with at most 2 queries.
A.3 Additional Experimental Results
A.3.1 Metadata Statistics
We provide metadata statistics characterizing the generated reports of each benchmarked method in Table 6, as well as statistics related to our evaluation metrics in Table 7.
| Report Length | Citations | Cost | |||||
| Chars | Words | Sentences | # Unique Refs | # Inline Citations | Latency (s) | Dollar Cost (USD) | |
@@ A.3.1 Metadata Statistics
| Human-written Exemplars | 4381 | 497 | 28 | 23 | 27 | N/A | N/A |
@@
| DeepResearcher (Llama-4) | 2573 | 319 | 35 | 8 | 7 | 31 | 0.00 |
| STORM (Llama-4) | 2766 | 381 | 31 | 18 | 21 | 162 | 0.00 |
| OpenScholar (Llama-4) | 3513 | 483 | 26 | 9 | 19 | 87 | 0.00 |
| Search Agents | |||||||
| Search Agent (Llama-4) | 1968 | 258 | 16 | 9 | 5 | 20 | 0.00 |
| Search Agent (GPT-4.1) | 3168 | 404 | 16 | 10 | 61 | 39 | 0.07 |
| Search Agent (o3) | 3844 | 501 | 24 | 11 | 16 | 263 | 0.15 |
| Search Agent (Claude) | 3977 | 499 | 27 | 13 | 8 | 147 | 1.36 |
| Search Agent (Gemini) | 2810 | 395 | 19 | 6 | 8 | 442 | 0.11 |
@@
| OpenAI DeepResearch | 6577 | 864 | 74 | 17 | 6 | 630 | 5.02 |
| DeepScholar Reference Pipeline | |||||||
| DeepScholar-ref (Llama-4) | 3499 | 360 | 53 | 19 | 35 | 313 | 0.00 |
| DeepScholar-ref (GPT-4.1) | 7863 | 735 | 115 | 20 | 89 | 234 | 1.66 |
| DeepScholar-ref (GPT-4.1, o3) | 5726 | 617 | 70 | 17 | 42 | 276 | 1.15 |
| DeepScholar-ref (GPT-4.1, Claude) | 5855 | 618 | 72 | 17 | 40 | 334 | 1.23 |
| DeepScholar-ref (GPT-4.1, Gemini) | 5623 | 570 | 86 | 24 | 63 | 349 | 1.29 |
| Avg. value over human-written exemplars | Relevant Metric | |
| # Important References from ArXiv.org | 11.47 | Ref. Cov. |
A.3.2 Results on DeepScholar-Nov-2025
To study how our benchmark behaves under both domain and temporal shifts in the live arXiv API, we instantiate a second benchmark slice, in addition to DeepScholar-June-2025 (Section 2), which is built from 63 Computer Science papers on arXiv to study domain coverage and robustness of DeepScholar-bench in our main results.
DeepScholar-Nov-2025 contains 200 queries sampled from more than 75 distinct arXiv subject areas, spanning Computer Science, Physics, Quantitative Biology, Economics, and Quantitative Finance. We evaluate a subset of high-performing systems to check whether our conclusions remain stable: DeepScholar-ref instantiated with Llama-4-Scout-17B-16E and with GPT-4.1+o3, and the Search Agent baseline instantiated with the same model configurations.
Table 8 reports the resulting scores across all seven metrics. Overall, we observe patterns that are consistent with our main results on DeepScholar-June-2025 (Table 2). For both DeepScholar-ref and the Search Agent baseline, the o3-based variants substantially outperform their Llama-4-Scout counterparts on Organization, Nugget Coverage, Relevance Rate, Citation Precision, and Claim Coverage, while Reference Coverage and Document Importance remain broadly low across baselines. These trends mirror the relative ranking and qualitative gaps seen in our main benchmark slice, suggesting that our conclusions from DeepScholar-June-2025 generalize beyond queries related to Computer Science.
We emphasize that DeepScholar-Bench is defined by an automated data-curation and evaluation pipeline rather than a single fixed dataset. Instantiating new slices such as DeepScholar-Nov-2025 only requires specifying a set of query papers and a date range; the pipeline then automatically constructs the corresponding benchmark and produces scores under the same evaluation protocol. This design allows practitioners to easily create additional domain- or time-specific evaluations while remaining comparable to our core results.
| Model | Org. | Nug. Cov. | Rel. Rate | Ref Cov. | Doc Imp. | Cite-P | Claim Cov. |
| DeepScholar-ref (Llama-4-Scout) | 0.120 | 0.358 | 0.395 | 0.072 | 0.061 | 0.178 | 0.581 |
| DeepScholar-ref (GPT-4.1 + o3) | 0.578 | 0.479 | 0.568 | 0.087 | 0.054 | 0.563 | 0.578 |
| Search Agent (Llama-4-Scout) | 0.108 | 0.252 | 0.179 | 0.034 | 0.124 | 0.189 | 0.314 |
| Search Agent (o3) | 0.608 | 0.480 | 0.623 | 0.078 | 0.763 | 0.475 | 0.452 |
A.3.3 Ablation Study: Understanding the Performance Impact of the Chosen Query
Our main experiments use the paper abstract as the query description (Section 3.2). A natural question is whether our conclusions depend on this particular choice of query formulation. To assess this, we run an ablation in which we replace the abstract with two alternative, realistic user queries: (i) a two-sentence summary of the paper’s key idea (Key Idea) and (ii) a single research question describing the paper’s main goal (RQ). For each paper, we prompt an LLM to convert the abstract into these two alternative query formulations. We then re-run the full benchmark for each query version and compute system-level scores for all metrics across our main baselines.
Table 9 reports Pearson correlations between system-level scores obtained under different query formulations (rows) for each metric (columns) with statistical significance testing based on a permutation-based paired test (). Overall, we observe very strong agreement across query types: correlations are typically above for Organization, Nugget Coverage, Reference Coverage, Document Importance, and Claim Coverage, and above for Coverage Relevance Rate in all cases. All correlations between Key Idea and RQ are statistically significant, indicating that more natural, user-facing query formulations yield highly consistent system rankings.
The main sensitivity to query formulation arises for Document Importance and, to a lesser extent, Citation Precision. The abstract vs. Key Idea and abstract vs. RQ correlations on Document Importance, as well as the abstract vs. Key Idea correlation on Citation Precision, are not statistically significant, despite having large magnitudes (e.g., for Document Importance). This suggests that these two metrics are most sensitive to how the query is phrased, likely because small changes in the query can shift which highly-cited papers are retrieved and cited. In contrast, the remaining metrics exhibit high and statistically significant correlations across all query pairs. Taken together, these results indicate that our benchmark conclusions are broadly robust to reasonable variations in query formulation, with only modest sensitivity in how document-level importance and citation precision are expressed across different query types.
| Query pair | Org. | Nug. | Cov. Rel. Rate | Ref Cov | Doc Imp | Cite-P | Claim Cov () |
| Abstract vs. Key Idea | 0.997∗ | 0.980∗ | 0.979∗ | 0.992∗ | 0.992 | 0.589 | 0.841∗ |
| Abstract vs. RQ | 0.988∗ | 0.979∗ | 0.772∗ | 0.951∗ | 0.707 | 0.766∗ | 0.877∗ |
| Key Idea vs. RQ | 0.991∗ | 0.983∗ | 0.836∗ | 0.958∗ | 0.766∗ | 0.942∗ | 0.981∗ |
A.3.4 Human Evaluation Details
To assess whether LLM-based metrics reflect human expert judgments, we conduct a human evaluation with 11 annotators, all Computer Science PhD students from four research universities across North America. In total, we collect over 300 human annotations aimed to assess the agreement between humans and LLMs in order to validate the LLM-based judges introduced by our automated evaluation for assessing knowledge synthesis and retrieval quality. We describe our setup in detail below.
Knowledge Synthesis.
For Organization & Coherency, we sample queries and show annotators the human-written related work section alongside a system-generated report. For each pair, annotators indicate whether they prefer the system report, prefer the human-written exemplar, or consider them similarly organized. These labels are used to evaluate the pairwise comparison outcomes underlying our Organization metric. For Nugget Coverage, we first generate information nuggets from the human-written related work section. Annotators are then shown individual nuggets and asked to judge whether each nugget is vital for understanding the paper, okay, or irrelevant. These nugget-importance labels determine which nuggets are treated as essential when computing nugget coverage scores.
Reference coverage and essential citations.
To ground our Reference Coverage metric in human judgments, we ask each annotator to select a high quality paper whose related work section they are comfortable with. For that paper, the annotator identifies at least six references they consider important (i.e., references that should appear in a good related work section) and at least six references they consider not important (i.e., references that could be omitted or substituted without harming the quality of the section). These labels form gold sets of important versus non-important references, which we use both to evaluate Reference Coverage and to validate our LLM-based importance labels. We then compare human majority labels to predictions from our LLM-judge over more than 130 blind reference-level annotations. The resulting confusion matrix (rows = human labels, columns = LLM predictions) is shown in Table 4.
A.3.5 Manual Validation of LLM-based Evaluation
@@
| Nugget Coverage | Nugget Coverage (Supported / Partially Supp. / Not Supp.) | |
| Retrieval Relevance Rate | Graded Relevance (0/1/2) | |
| Reference Coverage | Reference Importance (Not Imp./ Imp.) | |
| Document Importance | N/A | N/A |
| Citation Precision | Entailment (Entailed / Not Entailed) | |
| Claim Coverage | Entailment (Entailed / Not Entailed) |